In dry zone of a Blast Furnace (BF), following data are given: BF gas temperature = 1000oC, gas flow rate=4000 Nm3 /m2 /hr, gas density at NTP =1.5×10-3 g/cc, viscosity at 1000oC = 480×10-6 g/cm-s, particle diameter = 4 cm, average pressure = 1.6 atm. 

a.Calculate the pressure drop (mm of H2O/cm) for gas voidage 0.25. 

b.Calculate the percentage decrease in pressure drop when the gas voidage is increased from 0.25 to 0.35.

c.Calculate the percentage increase in pressure drop when the gas voidage is decreased to 0.20 from 0.25. 


1
Expert's answer
2021-02-09T03:34:05-0500

a) P = F/A

F = 4000

DENSITY = MASS/AREA

AREA = DENSITY/MASS

AREA = 1.5*10-3/480*10-6 = 3.125*10-12

P=F/A

P=4000/3.125*10-12

P= 1/781250000


b)15%


c)7.5%